When the line x+√3=√3 reflected by x axis what is the equation of reflected line

Grade:12

1 Answers

Vikas TU
6 years ago
Equation of ray of light:(3)^1/2 y = -x +3This line make 150 degree with +ve x – axis.So this makes 60 degree with normal.So reflected light make 30 degree with +ve x – axis.Equation of reflected light:x + (y - root3) = 1/root3 (x-10)y = x / root3 + 3